OpenStudy (anonymous):

at height of h meters you can see V kilometers to the horizon. These numbers are related by the equation V=3.5sqrt h. A person can see 278 km to the horizon from an airplane window. How high is the airplane?

OpenStudy (anonymous):

It's just \[278=3.5\sqrt{h}\] I hope you can solve it afterwards?

OpenStudy (anonymous):

yeah i have no idea lol thats why im here

OpenStudy (anonymous):

crap its 378 not 278

OpenStudy (anonymous):

\[278=\frac{7}{2}\sqrt{h}\] multiply both sides by 2 and divide by 7 \[\frac{556}{7}=\sqrt{h}\] and square it \[h=\frac{309136}{49}\approx6308m\approx6.3km\]